An investor buys a discount bond today that promises a face value payout of $4,500 in exactly four years. The investor holds it for one year and then decides to sell it in a secondary market to a different investor. If the interest rate is 6%, what price does he sell it for?

3,460.66
3,555.0
3,778.29
3,899.85
Answer #1

Answer: Option C: $3778.29

Price of Bond=P/(1+r)^n

P=Face value

r=interest rate

n=years to maturity

He sell the bond at a price=4500/(1+6%)^3=$3778.29